x^3+4x^2+7x+28=0
find possible rational roots:
find possible actual roots:

Answer by jim_thompson5910(35256)

Possible Rational Roots: List all the factors of 28

Possible Rational Roots: 1, 2, 4, 7, 14, 28, -1, -2, -4, -7, -14, -28

Note: make sure you list the negative factors as well

Another Note: This trick only works if the leading coefficient is 1 or -1.
------------------------------------------------------------------------

Actual Roots:

x^3+4x^2+7x+28=0

(x^3+4x^2)+(7x+28)=0

x^2(x+4)+7(x+4)=0

(x^2+7)(x+4) = 0

x^2+7 = 0 or x+4 = 0

x^2 = -7 or x = -4

x = sqrt(-7), x = -sqrt(-7), or x = -4

x = i*sqrt(7), x = -i*sqrt(7), or x = -4

The actual roots are x = i*sqrt(7), x = -i*sqrt(7), or x = -4

The first two are complex or imaginary roots. The third root is real and rational.

The only actual rational root is x = -4